OpenStudy (anonymous):

Write a linear system - Two numbers add together to give 14 and differ by 4

OpenStudy (zzr0ck3r):

x+y = 14 y-x = 4

OpenStudy (lgbasallote):

"two numbers add together to give 14" that would be x + y = 14 "differ by 4" mean x - y = 4 :)

OpenStudy (zzr0ck3r):

the difference in 2 and 4 is 2 thus 4-2 not 2-4

OpenStudy (zzr0ck3r):

i guess it does not matter here tho:)

OpenStudy (anonymous):

omg you guys are life savers ! Thank youuu